OpenStudy (hellokitty17):

Algebra 1 FLVS Help please!!! Will fan and MEDAL!!!

OpenStudy (hellokitty17):

The table below shows the distance y, in miles, traveled by a toy car in x minutes: Time (x) (minutes) 20 30 40 50 Distance (y) (miles) 5 10 15 20 Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and distance traveled by the toy car. (4 points) [Choose the value of correlation coefficient from 1, 0.8, 0.5, 0.02] Part B: What is the value of the slope of the graph of distance versus time, and what does the slope represent? (3 points) Part C: Does the data in the table represent correlation or causation? Explain your answer. (3 points)

OpenStudy (hellokitty17):

@imqwerty

OpenStudy (hellokitty17):

@Callisto @Dragonclaw88 @foreverthebeast @ganeshie8 @katt9421 @Jack1 @Mathie @NadiaElhaj_1

oregonduck (oregonduck):

First, plot the points. (20, 5), (30, 10), (40, 15), (50, 20) A. You'll see that it's a perfectly straight line, so it's a perfect positive correlation --- Positive because it increases from left to right. The value to represent this is 1. A correlation of 1 means time and distance have a perfect positive linear relationship, meaning if x increases by a certain value, y also increases by a certain value at the same time. In our case, every time x increases by 10, y increases by 5. B. Pick any two points and plug them into the slope formula. Lets use (20, 5) and (30, 10). m = (y₂ - y₁) / (x₂ - x₁)      = (10 - 5) / (30 - 20)      = 5/10      = 1/2 The slope represents the rate (or speed). Distance, rate, and time are related by              distance = rate • time distance / time = rate             rise / run = m                        1/2 = m The car's speed is 1/2 miles per minute. C. It's not causation because x must cause y to change, or vice versa. The car's distance has no effect on time. Time will increase steadily no matter what. Time has no effect on the car's distance. The car's movement is due to batteries, not because of time. It's a correlation because x and y are related by a common pattern mentioned in part A.

Nnesha (nnesha):

i don't get it. seriously

Nnesha (nnesha):

she is your sister right so why not just do it on a paper if you have to copy it from google ?

oregonduck (oregonduck):

yes so want me to do it that way

Nnesha (nnesha):

okay. sigh*

oregonduck (oregonduck):

plus we are in different houses and this is the only way i can help her

OpenStudy (hoothoot):

hey np dud i got u covered

oregonduck (oregonduck):

thxs man

OpenStudy (hoothoot):

i have a dude in collage well breaking from collage getting your answer

Nnesha (nnesha):

okay so explain it! let her do some work or teach her how to search on google

OpenStudy (hoothoot):

im a guy

OpenStudy (hoothoot):

give my bro some time

Nnesha (nnesha):

3 weeks enough or not ?

Nnesha (nnesha):

anyway good luck gys!! :) wish you all the best!

oregonduck (oregonduck):

thxs

Nnesha (nnesha):

cool that's how you found that pic ? i like it!

OpenStudy (hellokitty17):

thanks

Nnesha (nnesha):

Yw

OpenStudy (hoothoot):

ok my bro almost has your answer he is getting it finished soon

oregonduck (oregonduck):

ok

Nnesha (nnesha):

answer is already p there ? isn't it ??

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @OregonDuck First, plot the points. (20, 5), (30, 10), (40, 15), (50, 20) A. You'll see that it's a perfectly straight line, so it's a perfect positive correlation --- Positive because it increases from left to right. The value to represent this is 1. A correlation of 1 means time and distance have a perfect positive linear relationship, meaning if x increases by a certain value, y also increases by a certain value at the same time. In our case, every time x increases by 10, y increases by 5. B. Pick any two points and plug them into the slope formula. Lets use (20, 5) and (30, 10). m = (y₂ - y₁) / (x₂ - x₁)      = (10 - 5) / (30 - 20)      = 5/10      = 1/2 The slope represents the rate (or speed). Distance, rate, and time are related by              distance = rate • time distance / time = rate             rise / run = m                        1/2 = m The car's speed is 1/2 miles per minute. C. It's not causation because x must cause y to change, or vice versa. The car's distance has no effect on time. Time will increase steadily no matter what. Time has no effect on the car's distance. The car's movement is due to batteries, not because of time. It's a correlation because x and y are related by a common pattern mentioned in part A. \(\color{blue}{\text{End of Quote}}\) ths one ? and just a suggestion try to write it in ur own words otherwise you teacher will find out the original link :)

OpenStudy (hoothoot):

correlation coefficient is 1 slope is 5/10 which reduces to 1/2 not sure about the 3rd one